OpenStudy (xxsagexx):

Frank needs to fill juice cups for his little sister's birthday party. There are 35 juice cups and each juice cup holds 12 fluid ounces. The juice comes in 1-gallon bottles. How many 1-gallon bottles of juice will Frank need to purchase? (1 quart = 32 fluid ounces, 1 gallon = 4 quarts) (2 points) 1 bottle 2 bottles 3 bottles 4 bottles

OpenStudy (danjs):

\[\frac{ 35~cups }{ 1 }*\frac{ 12 ~oz }{ 1~cups }*\frac{ 1~quart }{ 32~oz }*\frac{ 1~gal }{ 4~quart }=\frac{ 35*12 ~~gal }{ 32*4 }\]

OpenStudy (danjs):

Round to the next highest whole number of gallons